chebysshev's theorem

chebysshev's theorem
Answer #1

Chebyshev’s Theorem:

at least 1-\frac{1}{k^2} of the data lie within k   standard deviations of the mean, that is, in the interval with endpoints \bar{x}-ks \ \ and \ \ \bar{x}+ks   for samples and with endpoints \mu-k\sigma \ \ and \ \ \mu+k\sigma for populations, where k is any positive whole number that is greater than 1.

For any numerical data set,

1.

  1. at least \frac{3}{4} of the data lie within two standard deviations of the mean, that is, in the interval with endpoints

\bar{x}-2s  and \bar{x}+2s for samples and with endpoints \mu-2\sigma and \mu+2\sigma for population

2.

at least  \frac{8}{9} of the data lie within three standard deviations of the mean, that is, in the interval with endpoints \bar{x}-3s   and \bar{x}+3s for samples and with endpoints \mu-3\sigma and \mu+3\sigma   for populations;